SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the May 14 machine background meeting

The meeting was limited to 1 hour duration because of the IR2 meeting scheduled at 9:30 am. This will be the case for all subsequent meetings until further notice. Note that the background group is expected to do a weekly reporting on Fridays at the Operational meeting at IR2, another good reason to stop at 9:30!

  1. Status of PEP-II and plans.
  2. During the owl May 14 shift, a 1 Hz injection HER beam was sent thru BABAR. No activity as sen on the PIN diodes. BABR took a run triggered on HER injection and was able to record beam activity in EMC and DIRC.

    The plan is to go and store both beams (actually performed on Sunday May 16) and go to collisions with B field=0 (foreseen now Wedneday). On Tuesday may 18, a higher current run is scheduled. (25 mA?)

  3. Early background measurements
  4. 2.1 Most tools are now in place : SVTRAD and analog system tested, signals available in EPICS and in MCC

    (see this recent posting :

    http://babar-hn.slac.stanford.edu:5090/HyperNews/get/babar_pepii/10.html)

    Still need to be done : more buttons file for Strip charts, correlation plots etc..

    1. Early look at DIRC signals caused by HER 1 Hz injection see
    2. http://babar-hn.slac.stanford.edu:5090/HyperNews/get/DetOper/439.html

      (Note 1.the plots are to be found in :

      http://www.slac.stanford.edu/~wormser/dirc_beam.ps

      Note 2. The 16 ns double peak structure in plot 3 has now been understood as a trigger jitter. Under investigation.)

    3. results were also reported by the EMC groupon the same data : low energy stuff (<20 MeV) in the forward region and in the horizontal plane.
  5. Experience from liaison shifts
  6. The general feeling is one of a good start. The points to improve are mainly :

    -Improve communication with BABAR/IR2. Telephoen is not always adequate. Suggestion is intercom + computer communication (talk, phone ,chat or any similar product)

    -Standardize logbook entries in BABAR logbook

    -refine to-do list, document correlation plots which are been taken ,etc..

    -Include liaison shifter names in BABAR shift list and in shift reports

    -Standardize shift and daily reporting at the 8 :00 meeting

    -Clarify operational procedures in case of problems with the beam abort modules. On that point, it was decied that the liaison person should refer this problem to the BABAR shift leader that will call the experts (the SVTRAD experts will be on shift permanently for the time being).

    These issues have been brought up by GW to the BABAR operational meeting in Ir2 at 9.30 teh same day and most of them have been implemented.

  7. Documentation

The two main documents were at that time

-The liaison shift manual

-The document for PEP-II operator

Please send comments updates, etc to Terry Geld and myself, respectively. Since the meeting, Thomas has also produced documentation on the Beam Abort module